Accelerator Update for September 3, 2008 - September 5, 2008
Wednesday Sept 3
The day shift began with Operations monitoring the Tevatron (TeV) store 6397 (74.32E30), the Recycler (RR) stash (327.8E10), and the Antiproton Source (Pbar) stack (32.0E10), with Pbar in standby due to a problem with a power supply for the Lithium Lens, with the Linac, Booster, and the Main Injector (MI) operating normally, with the Collider Detector at Fermilab (CDF) and DZero (D0) using the colliding beams, and with the MiniBooNE and NuMI experiments taking beam.
An RF expert brought a MI RF station (MIRF6) back online at 8:02 AM.
At 10 AM, BRF10 tripped off for the first of two times in five minutes. Operations contacted an RF expert after the second trip.
Operations terminated store 6397 at 10:24 AM.
Operations prepared Pbar for an access at 10:38 AM.
Operations put the TeV into shot setup at 10:39 AM.
Operations established store 6398 at 12:26 PM with an initial luminosity of 272.36E30.
A power supply expert reported 1:59 PM that he had repaired the lithium lens.
Pbar resumed stacking at 2:06 PM.
Operations reported at 3:44 PM that the next Meson MTest experiment (T978 – Calice) had received permission to take beam.
Operations reported at 3:58 PM that the Recycler had stashed antiprotons once during the day shift.
Operations began sending beam to Switchyard and MTest around 4:23 PM.
Operations disabled the MTest beam at 5:41 PM.
At 5:59 PM, a Booster vacuum pump (IPRF23) tripped off and caused BRF13 and 14 to also trip off. Everything reset.
At 6:01 PM, the P2 transfer line beam permit tripped off due to the sagging current on a power supply (QF12). Everything reset.
At 6:18 PM, the P2 transfer line tripped off due to QF12. It reset.
At 6:42 PM, the Pelletron VME tripped off. It recovered, but a vacuum pump in the Pelletron (R:IPA03I) tripped off and would not reset. Experts reported that this was not a serious problem.
At 7:23 PM, a Booster vacuum pump (IPRF17) tripped off and caused BRF7 and 8 to also trip. Everything reset.
At 11:30 PM, the P2 line tripped off due to QF12. This trip took down NuMI. Everything reset.
Operations reported at midnight that the Recycler had stashed six times during the evening shift.
Thursday Sept 4
The midnight shift began with Operations monitoring store 6398 (86.12E30) and the stash (197.4E10), with the Pbar stacking (23.0E10), and with the MiniBooNE and NuMI experiments taking beam.
At 12:42 AM, the P2 line tripped off due to QF12 for the first of four times in six hours. It reset.
At 12:59 AM, the NuMI beam permit tripped off due to a MI RF gap envelope problem. No MI RF stations had tripped. The beam permit reset.
At 1:02 AM, the NuMI beam permit tripped off again. This time MIRF6 had tripped off. Everything reset.
At 1:06 AM, BRF16 tripped off for the first of four times in three hours. It reset.
At 1:21 AM, NuMI tripped off again due to MIRF6.
At 2:46 AM, MIRF6 caused NuMI to trip off. Operators bypassed the station.
Operations established beam to MTest at 4:37 AM.
At 7:03 AM, Operations received a code one building alarm for the MCR. Operators turned off the Master Beam Switch and pulled the HEP key.
Operations reported at 7:13 AM that there had been no beam for about six minutes. Operators discovered that FESS personnel were conducting a pump test at Wilson hall. The test had caused a pressure alarm in the lower Linac Gallery.
Operations reported at 7:59 AM that the Recycler had stashed five times during the midnight shift.
Operations terminated store 6398 at 8 AM.
An RF expert reported at 8:06 AM that he had brought MIRF6 back online.
Operations put the TeV into shot setup at 8:33 AM.
Operations established store 6399 at 9:58 AM with an initial luminosity of 289.36E30.
A MI expert began a 90-minute study period at 10:27 AM.
A RF expert began work on MIRF6 at 1:36 PM.
A Linac expert began work on LRF4 at 1:54 PM.
MI resumed sending beam at 2:11 PM.
At 3:08 PM, the P2 line tripped off due to a correction magnet power supply (VTF12) trip. Everything reset.
Booster WAPS (West Anode Power Supply) tripped off at 5:00 PM. The trip was caused by rain water dripping onto BRF10, Experts are investigating.
At 5:51 PM, experts took BRF10 and BRF9 offline and added the spare BRF19. Booster is currently running at reduced intensity.
MTest turned off at 5:56 PM.
At 6:28 PM, BRF9 was brought back online, BRF10 remains offline for repairs. Booster returns to normal intensities.
Operations terminated store 6399 at 9:07 PM.
D0 went to controlled access at 9:28 PM, they finished at 10:16 PM.
Friday Sept 5
The midnight shift began with Operations monitoring the stash (288.8E10), with the Pbar stacking (22.9E10), and with the MiniBooNE and NuMI experiments taking beam.
Operations put the TeV into shot setup at 12:21 AM.
Operations established store 6403 at 2:39 AM with an initial luminosity of 225.1E30.
Operations established beam to MTest at 4:09 AM.
At 4:20 AM, LRF5 tripped off on permit inhibit for the first of many times over the next three hours. Experts are investigating.
At 5:00 AM, the PBar lens power supply (D:LNV) tripped off , investigating revealed that the bias power supply is bad, Experts were called in to investigate.
The Plans for Friday and the Weekend
The plans for today and the weekend are to repair the Pbar Lens, stack, stash, and store.
Comments and Suggestions
|last modified 09/05/2008 email Fermilab|
|Security, Privacy, Legal|